# Self-testing in the compiled setting via tilted-CHSH inequalities

### v1

@@ Abstract
In a Bell scenario, a classical verifier interacts with two non-communicating provers, producing a correlation. Certain correlations allow the verifier to certify, or self-test, the underlying quantum state and measurements. Notably, the family of tilted-CHSH inequalities has been used to self-test any two-qubit state. Self-tests underpin numerous device-independent protocols, however, a significant drawback of such applications is the non-communicating assumption. To address the non-communication assumption Kalai et al. (STOC ’23) give a procedure which compiles a bipartite Bell scenario into a 2-round interaction between a verifier and a single computationally bounded prover. In this work, we formalize self-testing for compiled Bell scenarios. We prove the maximal quantum value is preserved under compilation for the family of tilted-CHSH inequalities, and that any maximal violation constitutes a compiled self-test. More specifically, we establish the existence of an efficient isometry recovering the state and measurements in the second round.
@@ 1 Introduction
In a bipartite Bell scenario, two non-communicating provers receive inputs $x$ and $y$ and reply with outputs $a$ and $b$. The probabilities of $a,b$ given $x,y$ determine a correlation $\mathsf{p}={p(a,b|x,y)}$. Bell’s celebrated theorem implies that if the provers are permitted to share an entangled quantum state and make local quantum measurements, called a quantum model, then some behaviours have no realization under a local hidden-variable (LHV) model [Bel64]. The distinction between quantum and classical behaviours is often analyzed via Bell inequalities, which are linear inequalities on the behaviour satisfied by all LHV models, but can be violated by particular entangled quantum systems. The quintessential example is the CHSH inequality, considered in [CHSH69]. Bell inequality violations play a major role in areas of quantum computing and communication such as device-independent cryptography [ABG+07, Col07, PAM+10, PAB+09, Sca12], protocols for verifiable delegated quantum computation [RUV13, Gri19], and in the study of multiparty interactive proofs (MIPs) and the variant MIP with entangled provers [CHTW04, JNV+21].
Many key applications of Bell inequalities rely on a remarkable property known as self-testing [MY04, MYS12, YN13, ŠB20]. Informally, a Bell inequality is a self-test for an honest (ideal) quantum model $\mathrm{Q}$ if there exist local quantum operations which transform any optimal quantum model $\mathrm{Q}^{\prime}$ into the honest model $\mathrm{Q}$. It is well-known that the CHSH inequality is a self-test for a quantum model containing the maximally entangled state on two qubits [MY04]. Another notable example is the family of tilted-CHSH inequalities which have been used to self-test any bipartite state on two qubits [AMP12, YN13, BP15]. This latter family was instrumental in the celebrated result of Coladengelo, Tong Goh, and Scarani who employed these inequalities to self-test any pure bipartite state [CGS17].
Despite the enormous success of self-testing, a significant drawback in several applications is the assumption of non-communicating provers. Recently, a number of cryptographic approaches have been proposed that replace the non-communication assumption with computational assumptions [KLVY23, MV21, KMCVY22]. One approach, the KLVY compilation procedure introduced in [KLVY23], transforms a $2$-prover $1$-round Bell scenario, into a $1$-prover $2$-round scenario, with a computationally bounded prover. The core ingredient in the KLVY compilation procedure is quantum homomorphic encryption (QHE), which is used to simulate non-communication between the rounds of interaction. Essentially, the QHE scheme allows a quantum prover to perform honest measurements without knowing the given measurement settings.
@@
For the case of classical provers, it was shown that the KLVY compilation procedure preserves the winning probabilities [KLVY23]. In a follow-up work by Natarajan and Zhang [NZ23], it was shown that the maximal quantum violation of the CHSH inequality in the compiled setting is bounded by the maximal violation in the bipartite setting. Subsequent works have studied the KLVY compilation procedure in various contexts, including demonstrating quantum soundness for all 2-player XOR nonlocal games [CMM+24], Bell inequalities tailored to maximally entangled bipartite states [BVB+24], delegated quantum computation with a single-device [NZ23, MNZ24], and even in the study of contextuality [ABCC24]. Despite these advancements, many foundational results in the bipartite setting, such as the self-testing for all bipartite qubit states  [YN13, BP15] have yet to be reproduced in the compiled setting.
Summary of results
We show that for a family of extended tilted-CHSH inequalities [BSB24], the maximal quantum violation in the compiled setting is bounded by the maximal quantum violation in the bipartite setting. To obtain this bound, we establish a technical result which extends the powerful sum-of-squares technique to allow terms with arbitrarily high degrees in Bob’s operators, partially answering a question posed in [CMM+24]. Our next contribution is a step towards a mathematical theory of self-testing in the compiled setting. We introduce a notion of self-testing in the compiled setting, which certifies the second-round measurements and the post-measurement state after the first round of interaction. Lastly, we combine these points by showing that the extended tilted-CHSH inequalities are self-tests according to our definition. In summary:
  • In Lemma 3.2 we extend the SOS proof techniques of Cui et al. [CMM+24] to allow for monomials with arbitrarily high degree in Bob operators.
  • In Theorem 4.2 we prove that the quantum bound for an extended family of tilted-CHSH expressions is preserved under the KLVY compilation procedure.
  • In Definition 3.4 we introduce a notion of self-testing for compiled Bell inequalities, which certifies the structure of the measurements in the second round and the state of the device after the first round of the protocol.
  • In Theorem 4.4 we combine the previous two points to show that the extended tilted-CHSH inequalities are a self-test according to our definition.
1.1 Technical discussion
In this section, we briefly discuss the KLVY compilation procedure and the family of extended tilted-CHSH inequalities. Additionally, we introduce the compiled counterpart to a bipartite quantum model, which is essential for our approach to self-testing. Lastly, we review self-testing in the bipartite setting and give an informal version of our definition of self-testing in the compiled setting.
The KLVY Compiler
The KLVY compiler transforms any Bell scenario into an interaction between a verifier and a single computationally-bounded prover, who sequentially assumes the roles of the two provers Alice and Bob. The procedure is roughly as follows:
  1. 1.
    The prover prepares an initial state $|\psi\rangle$.
  2. 2.
    The verifier samples challenges $(x,y)$, sends the encrypted first challenge $\chi=\mathsf{Enc}(x)$ to the prover, who acts as Alice, measures the state, and returns the outcome $\alpha$.
  3. 3.
    The verifier then sends $y$ in the clear, and the prover, acting as Bob, measures the residual state and returns the outcome $b$.
  4. 4.
    The verifier determines the decrypted answer pair $(a=\mathsf{Dec}(\alpha),b)$.
The compiled value of a Bell inequality is always at least the quantum value, as any quantum strategy can be implemented with homomorphic encryption [KLVY23]. Establishing upper bounds on the compiled value is challenging, since general techniques for bounding the value of a nonlocal game depend on the spatial separation between the two provers.
Extended tilted-CHSH inequalities
@@ 1.1 Technical discussion
The family of “tilted-CHSH” inequalities was introduced by Acín, Massar, and Pironio [AMP12]. The Bell operators for these inequalities take the form $\alpha_{\theta}\langle A_{0}\rangle+\langle A_{0}B_{0}\rangle+\langle A_{0}B_{1}\rangle+\langle A_{1}B_{0}\rangle-\langle A_{1}B_{1}\rangle$, where $\langle A_{x}B_{y}\rangle,\langle A_{x}\rangle$ denote the expectation of measurements corresponding to settings $X=x,Y=y$, and $\alpha_{\theta}\in\mathbb{R}$. This family of inequalities has applications in self-testing, notably, they are tailored to robustly self-test the 2-qubit states $\cos(\theta)|00\rangle+\sin(\theta)|11\rangle$ [YN13, BP15]. Subsequently, these inequalities were used as part of a more complex protocol to obtain self-testing for all pure bipartite quantum states [CGS17]. The work of Barizien, Sekatski, and Bancal [BSB24] extended this family to include extra degrees of freedom in Bob’s measurements, which we will refer to as “extended” tilted-CHSH inequalities.
Compiled counterpart of a bipartite model
A bipartite (quantum) model $\mathrm{Q}$, consists of a shared state $|\phi\rangle$, along with POVM measurements $\{P_{a|x}\}$ and $\{Q_{b|y}\}$ for Alice and Bob, respectively. In the bipartite setting, one can consider the situation where Alice measures first using POVM $\{P_{a|x}\}$, collapsing the state to a post-measurement state $\rho_{a|x}$ on Bob’s space, upon which Bob performs his measurement $\{Q_{b|y}\}$. In the case that $\rho_{a|x}$ is pure, we write $\rho_{a|x}=|\phi_{a|x}\rangle\!\langle\phi_{a|x}|$. Analogously, we consider a compiled (quantum) model $\widetilde{\mathrm{Q}}$ consisting of a family of post-measurement states for “Alice” $|\widetilde{\phi}_{\alpha|\chi}\rangle$, which correspond to the state of the device following the encrypted question $\chi$, and encrypted answer $\alpha$, and a POVM $\{N_{b|y}\}$ employed by “Bob”. One could also consider a more general compiled quantum model which includes a description of the initial state and Alice’s operators. The point of taking the coarser model is that it allows us to introduce the notion of the compiled counterpart of a bipartite model $\mathrm{Q}$, which relates the post-measurement information in bipartite setting with that of the compiled setting.
Definition 1.1 (Compiled counterpart (informal)).
Given a bipartite model $\mathrm{Q}$ for which $\rho_{a|x}$ is pure, the compiled counter-part $\widetilde{\mathrm{Q}}|_{B}$, is a compiled model $\widetilde{\mathrm{Q}}|_{B}=(\{|\widetilde{\phi}_{\alpha|\chi}\rangle\},\{\{N_{b|y}\}_{b\in\mathcal{B}}\}$) satisfying the following conditions:
$|\widetilde{\phi}_{\alpha|\chi}\rangle$ $=|\phi_{a|x}\rangle,\ \text{ for all }\chi:\mathsf{Enc}(x)=\chi,\ \alpha:\mathsf{Dec}(\alpha)=a.$
$N_{b|y}$ $=Q_{b|y},\ \text{ for all }b,y.$
Strictly speaking, Definition 1.1 is not a full counter-part of a quantum model $\mathrm{Q}$, but is more accurately described as a counter-part to Bob’s view of $\mathrm{Q}$. We also note that this definition requires that the states $\rho_{a|x}$, which are obtained by applying Alice’s measurement and tracing out, are pure. This condition will be satisfied when we take $\mathrm{Q}$ to be the honest quantum model for the tilted-CHSH inequalities.
Self-testing in the Bipartite Setting
A Bell inequality, $I\leq\eta^{Q}$, self-tests an optimal, also called honest, bipartite model $\mathrm{Q}^{*}$, if any optimal bipartite model is essentially the same as $\mathrm{Q}^{*}$, modulo, some physically irrelevant degrees of freedom. This is more formally treated by the existence of local isometries which maps the observed model to the honest one.
Definition 1.2 (Bipartite self-test (informal)).
A Bell inequality, $I\leq\eta$, is a self-test for a bipartite model $\mathrm{Q}^{*}=\left(\{P_{a|x}\},\{Q_{b|y}\},|\phi\rangle\right)$ if, for any bipartite model $\mathrm{Q}=\left(\{M_{a|x}\},\{N_{b|y}\},|\Psi\rangle\right)$ that saturates the bound $\eta$, there exist an auxiliary state $|\zeta\rangle$ and local isometries $V_{A}$ and $V_{B}$, that satisfy the following condition for all $x,y,a,b$:
$V_{A}\otimes V_{B}(M_{a|x}\otimes N_{b|y})|\Psi\rangle=(P_{a|x}\otimes Q_{b|y})|\phi\rangle\otimes|\zeta\rangle,$
Self-testing in the Compiled Setting
There are two difficulties in extending the notion of self-testing to the compiled setting. First, what are the relevant or irrelevant degrees of freedom in the compiled setting? Indeed, if we consider arbitrary isometries, then it is always possible to extract the ideal model $\widetilde{\mathrm{Q}}^{*}$. Secondly, using current techniques it is unclear how to make determinations about Alice’s operators. In particular, rigidity relations enjoyed by Alice’s operators in the bipartite model, such as anti-commutation, may not hold for the measurements performed by the corresponding homomorphic evaluation. Below we outline a notion of self-testing that naturally addresses the first concern, by requiring the isometry to be efficient, and somewhat sidesteps the second concern by focusing on the post-measurement perspective after which Alice operators have already been applied. We note that our definition covers self-testing type results from [NZ23, CMM+24, BVB+24].
Definition 1.3 (Compiled self-test (informal)).
Let $\mathrm{Q}^{*}=(\{P_{a|x}\},\{Q_{b|y}\},|\phi\rangle)$, denote the honest bipartite model. A Bell inequality $I\leq\eta$ is a self-test for the corresponding compiled counter-part $\widetilde{\mathrm{Q}}^{*}|_{B}=(\{|\widetilde{\phi}_{\alpha|\chi}\rangle\},\{Q_{b|y}\})$, if for every complied model $\widetilde{\mathrm{Q}}=\left(\{|\Psi_{\alpha|\chi}\rangle\},\{N_{b|y}\}\right)$ that achieves $I=\eta$, there exists a negligible function $\delta_{\mathsf{QHE}}$, an efficient isometry $V$, and auxiliary states $|\mathsf{aux}_{\alpha|\chi}\rangle$, which satisfy the following for all $x,b,y$:
$\mathop{\mathbb{E}}_{\chi:\mathsf{Enc}(x)=\chi}\sum_{\alpha}\big{\|}VN_{b|y}|\Psi_{\alpha|\chi}\rangle-Q_{b|y}|\widetilde{\phi}_{\alpha|\chi}\rangle\otimes|\mathsf{aux}_{\alpha|\chi}\rangle\big{\|}^{2}\leq O(\delta_{\mathsf{QHE}}).$
1.2 Outline of proof techniques
Upper Bounding the Complied Value
Upper bounds on the compiled quantum value of certain Bell inequality can often be verified using the well-known sum-of-squares (SOS) technique [NZ23, CMM+24, BVB+24]. Informally, this technique [^1] We suggest reading Section 4 of [CMM+24] for a more formal discussion of this technique. relates the maximum compiled value $\eta$, of a Bell functional $I$, to a suitably nice decomposition of the Bell operator or Bell polynomial $S$ as a sum of hermitian squares,
$\eta\mathbbm{1}-S=\sum_{i}P^{\dagger}_{i}P_{i}.$
Here, suitably nice requires that the terms $P_{i}$ involved in the decomposition are degree $2$ monomials in Alice and Bob’s observables.
Our first technical result provides an extension of the aforementioned SOS technique to a larger family of Bell polynomials. More specifically, we extend the pseudo-expectation techniques in [NZ23, CMM+24] to allow for evaluations on polynomial terms $P_{i}$ that consist of arbitrary monomials in the algebra generated by Bob’s observables. In Lemma 3.2 we prove that an extended pseudo-expectation will be positive on the corresponding hermitian square $P_{i}^{\dagger}P_{i}$ for any such term $P_{i}$. As a consequence of Lemma 3.2, we obtain the following general result that can accommodate SOS decompositions of higher degree in Bob’s observables [^2] Although our set-up and proof of Lemma 3.2 are given for binary observables, we expect this approach can generalize to larger question sets..
Theorem 1.4 (Informal).
Let $\mathcal{S}$ be a Bell scenario and let $S$ be a Bell polynomial. Suppose that $\eta\mathbbm{1}-S=\sum P^{\dagger}_{i}P_{i}$ is a SOS decomposition with each term $P_{i}$ of the form
$P_{i}=\sum_{j}\gamma_{j}(A_{x})^{k_{j}}w_{j}(B),$ (1.1)
for some $\gamma_{j}\in\mathbb{C}$, $k_{j}\in\{0,1\}$ and, $w_{j}(B)$, an arbitrary monomial in Bob’s observables. Then $\eta$ is an upper bound on the compiled value of $S$.
@@ 1.2 Outline of proof techniques
Next, we apply our general result to the SOS decomposition for the extended tilted-CHSH inequalities presented in [BSB24]. In Theorem 4.2 we use our new result to prove that the maximum quantum value achieved for any of the extended tilted-CHSH functionals is preserved by the KLVY compilation procedure.
Compiled extended tilted-CHSH inequalities are compiled partial self-tests
In the compiled setting, we show that our definition of self-testing is satisfied by the extended family of tilted-CHSH inequalities. Our proof approach is inspired by the self-testing results of Bamps and Pironio [BP15] which show a near-optimal Bell violation in the compiled setting implies key relations on the expectation of each SOS polynomial. A key difference to the bipartite case is that, instead of proving relations of Alice’s and Bob’s operators on a shared bipartite state. The relations we prove capture the structure of the observables in the second round measured on the post-measurement state after the first round. Interestingly, these relations strongly resemble those found in the bipartite setting, indicating that our notion of a compiled self-test allows existing techniques to be applied.
@@
  • One central goal in the compiled setting is to determine if the KLVY compiler always preserves the bipartite quantum value. Bell inequalities offer more flexibility than nonlocal games, perhaps this could be utilized to find compiled Bell inequalities with greater quantum violations in the compiled case than in the bipartite case.
  • Tilted-CHSH inequalities were an integral component of the self-testing for all pure bipartite states [CGS17]. Building off of our work on compiled tilted-CHSH inequalities, a natural question is whether similar results can be obtained in the compiled setting.
  • It would be desirable to understand the fundamental limitations of our notion of self-testing and other similar notions such as the computational self-testing given in [MV21]? Furthermore, is a finer notion of self-testing in the compiled setting that characterizes both Alice and Bob’s operators and the initial state possible without specifying the underlying QHE scheme?
  • All current techniques for bounding the value of compiled nonlocal games/Bell inequalities can be obtained using some variant of the sum-of-squares decomposition approach. Given our improvements to this approach outlined in Lemma 3.2, it is possible to search for valid decompositions which include arbitrary words in Bob’s operators. Is it possible to use this approach to give a limited variant of the NPA hierarchy [NPA08] in the compiled setting?
Outline of the paper
The paper proceeds as follows. Section 2 covers our mathematical notation, preliminaries on bipartite Bell inequalities, definitions for efficient quantum operations, and a brief overview of quantum homomorphic encryption. In Section 3, we define compiled Bell inequalities for bounded quantum provers, discuss bounding the quantum value of a compiled Bell inequality, and define a notion of self-testing for compiled quantum models. In Section 4 we consider the family of compiled tilted-CHSH inequalities, and show that maximal Bell inequality violation in the compiled setting is bounded by the maximal violation in the bipartite setting. We also show that this family of compiled Bell inequalities satisfies our definition of compiled self-tests outlined in Section 3.
Acknowledgements
The authors thank Simon Schmidt, Ivan Šupić, Anand Natarajan and Tina Zhang for helpful discussions. AM is supported by NSERC Alliance Consortia Quantum grants, reference number: ALLRP 578455 - 22 and the NSERC Discovery Grants Program. CP acknowledges funding support under the Digital Horizon Europe project FoQaCiA, Foundations of quantum computational advantage, grant no. 101070558, together with the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). LW acknowledges funding support from the EPSRC grant EP/SO23607/1.
2 Preliminaries

### v3

@@ Abstract
This work investigates the family of extended tilted-CHSH inequalities in the single-prover cryptographic compiled setting. In particular, we show that a quantum polynomial-time prover can violate these Bell inequalities by at most negligibly more than the violation achieved by two non-communicating quantum provers. To obtain this result, we extend a sum-of-squares technique to monomials with arbitrarily high degree in the Bob operators and degree at most one in the Alice operators. We also introduce a notion of partial self-testing for the compiled setting, which resembles a weaker form of self-testing in the bipartite setting. As opposed to certifying the full model, partial self-testing attempts to certify the reduced states and measurements on separate subsystems. In the compiled setting, this is akin to the states after the first round of interaction and measurements made on that state. Lastly, we show that the extended tilted-CHSH inequalities satisfy this notion of a compiled self-test.
@@ 1 Introduction
In a bipartite Bell scenario, two non-communicating provers receive inputs $x$ and $y$ and reply with outputs $a$ and $b$ to a verifier. The collection of probabilities of observing outcomes $(a,b)$ given $(x,y)$ determines a correlation $\mathsf{p}=\{p(a,b|x,y)\}$. Bell’s celebrated theorem implies that if the provers are permitted to share an entangled quantum state and make local quantum measurements, called a bipartite (quantum) model, then certain correlations have no realization by a classical (or local hidden-variable) model [Bel64]. The distinction between quantum and classical correlations is often explored through Bell inequalities. A Bell inequality is a linear inequality on the set of correlations which is satisfied by all classical correlations. Hence, these inequalities can be violated by certain models using quantum entanglement, realizing correlations that are not classical. The quantum value of a Bell inequality refers to the largest violation achievable by a bipartite (quantum) model. A prominent example is the Clauser-Horne-Shimony-Holt (CHSH) inequality, where the classical bound is $2$, but the quantum value is $2\sqrt{2}$ [CHSH69].
Due to their ability to witness these non-classical effects, Bell inequality violations play a major role in areas like device-independent cryptography [ABG+07, Col07, PAM+10, PAB+09, Sca12], protocols for verifiable delegated quantum computation [RUV13, Gri19, CGJV24], and in the study of multiprover interactive proofs (MIPs) and the variant MIP with entangled provers [CHTW04], also called nonlocal games. Many of the key applications of Bell inequalities rely on a remarkable property known as self-testing [MY04, MYS12, YN13, ŠB20]. Informally, a Bell inequality is a self-test for an ideal bipartite (quantum) model $\mathrm{Q}$ if there exist local isometries which transform any employed bipartite model $\mathrm{Q}^{\prime}$ achieving maximum Bell violation into the ideal model $\mathrm{Q}$. It is well-known that the CHSH inequality is a self-test for the bipartite model employing a maximally entangled state on two qubits, along with the Pauli $\sigma_{x}$ and $\sigma_{z}$ measurements, among others [MY04]. Another prominent example is the family of tilted-CHSH inequalities [AMP12, YN13, BP15], which self-test partially entangled two-qubit states, and were integral in the work of Coladangelo, Goh, and Scarani who employed them as part of a protocol to self-test any pure bipartite entangled state [CGS17].
Despite the enormous success of self-testing, a practical drawback is the requirement of multiple non-communicating quantum provers. Recently, a number of cryptographic approaches have been proposed that replace the non-communication assumption with computational assumptions [KLVY23, MV21, KMCVY22]. This makes the setting more practical by having a single quantum prover, rather than multiple. One new and prominent approach is the Kalai-Lombardi-Vaikuntanathan-Yang (KLVY) compilation procedure introduced in [KLVY23], which transforms a $2$-prover 11-round Bell scenario into a $1$-prover 22-round scenario with a single computationally bounded prover. The core ingredient in the KLVY compilation procedure is quantum homomorphic encryption (QHE), which emulates, to a certain extent, the non-communication between the rounds of interaction. In the compiled game, the inputs to the prover happens sequentially. In the first round, the prover obtains an encryption $\chi$ of the input $x$ from the verifier. Without breaking the security, the prover cannot distinguish between encryptions of different inputs. The prover performs a polynomial time quantum circuit on $\chi$, and then returns an output $\alpha$ to the verifier. In the second round, the information about $x$ has already been “hidden” from the prover, so the verifier can send input $y$ in the plain (i.e. unencrypted) to the prover, upon which the prover can perform a measurement and return outcome $b$ to the verifier. The verifier checks for a Bell inequality violation (across many such interactions) using the values of $x$, the decryption of $\alpha$, along with $(y,b)$. QHE has two key features that makes this resemble the bipartite setting. Firstly, it allows the first round quantum prover to perform measurements as they would have in the bipartite setting, without knowing the input. Secondly, the encryption ensures that no classical polynomial-time prover can violate a Bell inequality by more than an negligible amount (see Section 3 details). Both of these facts are non-trivial and were the subject of [KLVY23].
@@
In a follow-up work, Natarajan and Zhang showed that the maximal quantum violation of the CHSH inequality in the compiled setting is bounded by the maximal violation in the bipartite setting, up to negligible factors in the security parameter [NZ23]. Subsequent works have analyzed the quantum soundness of the KLVY compilation procedure for other multiprover scenarious, including all 2-player XOR nonlocal games [CMM+24], Bell inequalities tailored to maximally entangled bipartite states [BVB+24], delegated quantum computation with a single-device [NZ23, MNZ24], and even in the study of contextuality [ABCC24]. Despite these advancements, many results have yet to be reproduced in the compiled setting. Our work takes another step in growing the list of protocols that will function as desired in the compiled setting.
Upper bounding compiled Bell violations
As mentioned, the compiled value of a Bell inequality is always at least the quantum value. This is because any bipartite (quantum) model can be implemented with homomorphic encryption via a correctness property of the QHE scheme used in the procedure. On the other hand, establishing upper bounds on the largest violation possible in the compiled setting is challenging, as general techniques for bounding these violations depend on the spatial separation between the two provers. Nonetheless, upper bounds on the violations of a certain Bell inequalities in the compiled setting can be verified using the sum-of-squares (SOS) technique [NZ23, CMM+24, BVB+24]. The SOS approach is a powerful method and has been used extensively to upper bound Bell inequality violations and the values of nonlocal games in the bipartite setting. Informally, this technique relates the maximum compiled value $\eta$, of a Bell functional $I$, to a decomposition of the Bell operator or Bell polynomial $S$ as a sum of Hermitian squares, $\eta\mathbbm{1}-S=\sum_{i}P^{\dagger}_{i}P_{i}$. Before our work, progress was made on realizing this approach in the compiled setting, however, there were some limitations. In particular, it was required that the polynomials $P_{i}$ involved in the decomposition were at most degree two in both Alice’s and Bob’s observables, restricting the technique to Bell inequalities with an SOS decomposition of this form; this excludes, for example, the family of tilted-CHSH inequalities.
Our first result extends the SOS technique to a larger family of Bell polynomials. More specifically, we extend the pseudo-expectation techniques in [NZ23, CMM+24] to allow for evaluations on polynomial terms $P_{i}$ that consist of arbitrary monomials in the algebra generated by Bob’s observables. In Theorem 3.3 we prove that an extended pseudo-expectation will be positive on the corresponding Hermitian square $P_{i}^{\dagger}P_{i}$ for any such term $P_{i}$. Consequently, we show that for any Bell inequality with an SOS decomposition in which $P_{i}$ are of the form $P_{i}=\sum_{j}\gamma_{j}(A_{x})^{k_{j}}w_{j}(B)$ for some $\gamma_{j}\in\mathbb{C}$, $k_{j}\in\{0,1\}$ and $w_{j}(B)$ being arbitrary monomials in Bob’s observables, $\eta$ is an upper-bound on the maximum compiled quantum value. Our extension captures a wide class of Bell inequalities including tilted-CHSH, enabling us to bound the compiled value of the tilted-CHSH inequalities, by the quantum value and a negligible function of security parameter, see Theorem 3.5 for details.
A compiled self-testing result
Our second contribution is a concept of self-testing in the compiled setting. One of the main obstacles to deriving self-testing results in the compiled setting is the lack of techniques for extracting any algebraic relations on the measurement operators acting under the encryption. Nevertheless, it remains possible to derive relations on the observables in the second round. With this in mind, we consider a partial notion of self-testing that applies to the measurements made by the prover in the second round. In particular, our definition only requires the existence of an isometry robustly certifying the ideal post-measurement state after the first round, and the action of the measurements made in the second; see Definition 4.5 for details.
@@
As our final result, we provide an example by showing violations of the compiled tilted-CHSH inequalities satisfy this notion of partial self-testing. This family of inequalities was introduced by Acín, Massar, and Pironio [AMP12], and the Bell functionals take the form $\alpha_{\theta}\langle A_{0}\rangle+\langle A_{0}B_{0}\rangle+\langle A_{0}B_{1}\rangle+\langle A_{1}B_{0}\rangle-\langle A_{1}B_{1}\rangle$, where $\langle A_{x}B_{y}\rangle,\ \langle A_{x}\rangle$ denote the expectation of measurements corresponding to settings X=x,Y=yX=x,\ Y=y, and $\alpha_{\theta}\in\mathbb{R}$. Notably, they are tailored to robustly self-test the two qubit states $\cos(\theta)|00\rangle+\sin(\theta)|11\rangle$ [YN13, BP15], and were used as part of a more complex protocol to obtain self-testing for all pure bipartite entangled states [CGS17]. The work of Barizien, Sekatski, and Bancal [BSB24] extended this family to include extra degrees of freedom in Bob’s measurements, which we will refer to as “extended” tilted-CHSH inequalities.
@@
We apply Theorem 3.3 to the SOS decomposition for the extended tilted-CHSH inequalities presented in [BSB24]. Specifically, in Theorem 3.5 we prove that the maximum quantum value achieved for any of the extended tilted-CHSH functionals is preserved by the KLVY compilation procedure. Then in Theorem 4.6 we use this same decomposition to prove that this family of games is a compiled self-test according to Definition 4.5.
Related work
A recent work of [KMP+24] implies that the compiled value of any 2-prover Bell scenario is bounded by the largest violation possible among so-called commuting operator models. However, unlike some previous results, such as [BVB+24, CMM+24], the upper bound in [KMP+24] lacks a dependence on the security parameter $\lambda$, making it unclear how the compiled value is related to the quantum value at fixed security parameters. Hence, results such as ours, which obtain a bound on the compiled value that depends negligibly on the security parameter, remain of great importance. Furthermore, [KMP+24] also considers a notion of self-testing in the compiled setting, however, due to their methods the results are in terms of commuting operator self-tests (as defined in [PSZZ23, Proposition 7.8]) and only hold in the limit of the security parameter $\lambda\rightarrow\infty$.
Another related work is [MV21], which presents a protocol for certifying that an unknown computationally bounded device has prepared a maximally entangled pair of qubits, and whether a measurement was performed on each qubit in either the computational or Hadamard basis. The techniques used to prove our compiled self-test have similarities to those of [MV21], particularly in the choice of isometry (see Definition 4.2) and proof structure, which in turn resembles self-testing techniques in the bipartite setting [BP15]. There are however some key differences. Firstly, [MV21] certifies the preparation of a maximally entangled state by the device before any measurements are made. While our results are tailored to the more general class of partially entangled states, we only make statements about the post-measurement states after each round. It is an interesting open question if our results can be extended in this way (see Section 4.1 for more details), and statements weaker than certifying the prepared state could also be possible. For example, can a compiled self-test be used to show the prepared state must have been entangled? Another significant difference to [MV21] is that the self-testing protocol in this work strongly resembles the bipartite case, owing to the compilation procedure mapping bipartite nonlocal scenarios to single prover scenarios. Our main result can therefore be interpreted as translating a self-testing statement in the Bell scenario to one in the compiled Bell scenario. On the other hand, the authors of [MV21] describe their approach as more “custom”, guided by the available cryptographic primitives, and pose the open question of finding a general procedure for translating self-testing results from the nonlocal setting. We showed this is possible for the special case of titled-CHSH inequalities.
@@
  1. 1.
    Tilted-CHSH inequalities were an integral component of the self-testing for all pure bipartite entangled states [CGS17]. Building off of our work on compiled tilted-CHSH inequalities, a natural question is whether similar results can be obtained in the compiled setting.
  2. 2.
    It would be desirable to understand the fundamental limitations of our notion of self-testing and other similar notions such as the computational self-testing given in [MV21]. Furthermore, is a finer notion of self-testing in the compiled setting that characterizes both Alice’s and Bob’s operators and the initial state possible without specifying the underlying QHE scheme? Moreover, is every self-test in the standard Bell scenario also a compiled self-test, and vice-versa?
  3. 3.
    Many current techniques for bounding the value of compiled nonlocal games/Bell inequalities can be obtained using some variant of the sum-of-squares decomposition approach. Given our improvements to this approach outlined in Theorem 3.3, it is possible to search for valid decompositions which include arbitrary words in Bob’s operators. Is it possible to use this approach to give a limited variant of the NPA hierarchy [NPA08] in the compiled setting?
Acknowledgments
The authors thank Simon Schmidt, Ivan Šupić, Anand Natarajan, and Tina Zhang for helpful discussions. We also thank the anonymous TQC reviewers for valuable feedback. AM is supported by the NSERC Alliance Consortia Quantum grants program, reference number: ALLRP 578455 - 22 and the NSERC Discovery Grants program 2024-06049. CP is supported by the Digital Horizon Europe project FoQaCiA, Foundations of quantum computational advantage, grant no. 101070558, and the Natural Sciences and Engineering Research Council of Canada (NSERC) reference number: ALLRP 578455 - 22. LW is supported by the Engineering and Physical Sciences Research Council (EPSRC Grant No. EP/SO23607/1) and the European Union’s Horizon Europe research and innovation programme under the project “Quantum Security Networks Partnership” (QSNP, grant agreement No. 101114043).
2 Background